OpenStudy (shelld93):

In a certain university course 8 out of every 10 students finish it. If in this semester 94 students are in the course, how many are expected to finish ?

OpenStudy (accessdenied):

For 10 students attending, 8 students finish. If we assume the ratio is expected for any amount of attendees: If 10 students attend, then 8 students succeed. If 94 students attend, then x students succeed. Do you know how to write this proportion?

OpenStudy (shelld93):

no. it kinda seems like youve taken my question and just re worded it :/ im not that bright so sorry if im a bit slow.

OpenStudy (accessdenied):

Yep, that's essentially what I did. And its fine. I reworded it so that the information lines up a bit better. The ratio is just a division: \( \dfrac{8 \text{ students finish}}{10\text{ students total}} \) , \( \dfrac{ x \text{ students finish}}{94 \text{ students total}} \) These would be equal. Removing the words, we just have the proportion which we may solve. \( \dfrac{x}{94} = \dfrac{8}{10} \)

OpenStudy (shelld93):

fractions. yay :/ okay so where do i go from here ?

OpenStudy (anonymous):

Do you know how to solve a proportion?

OpenStudy (anonymous):

I take that as a no?

OpenStudy (shelld93):

no i dont :/

OpenStudy (accessdenied):

I think the phrase "cross-multiply and divide" is often associated with proportions, if that rings a bell.

OpenStudy (shelld93):

sort of.

OpenStudy (anonymous):

Cross multiply.\[94\times8=?\]Divide\[?\div10=?\]

OpenStudy (shelld93):

752 752 divided by 10 = 75.2

OpenStudy (anonymous):

That's what I got. :)

OpenStudy (shelld93):

so 75 students are expected to finish the course. cant have 75.2 people. and because th is below 5 i round down yeah ?

OpenStudy (accessdenied):

Yep, that sounds reasonable. Can't really have a fifth of a person finish. :)

OpenStudy (anonymous):

That's what I would do.

OpenStudy (shelld93):

thanks heaps !!!!! i really appreciate it !

OpenStudy (accessdenied):

Youŕe welcome!